# Prompting LLMs for Code Editing: Struggles and Remedies

### v1

@@ Abstract.
Large Language Models (LLMs) are rapidly transforming software engineering, with coding assistants embedded in an IDE becoming increasingly prevalent. While research has focused on improving the tools and understanding developer perceptions, a critical gap exists in understanding how developers actually use these tools in their daily workflows, and, crucially, where they struggle. This paper addresses part of this gap through a multi-phased investigation of developer interactions with an LLM-powered code editing and transformation feature—Transform Code—in an IDE widely used at Google. First, we analyze telemetry logs of the feature usage, revealing that frequent re-prompting can be an indicator of developer struggles with using Transform Code. Second, we conduct a qualitative analysis of unsatisfactory requests, identifying five key categories of information often missing from developer prompts. Finally, based on these findings, we propose and evaluate a tool—AutoPrompter—for automatically improving prompts by inferring missing information from the surrounding code context, leading to a 27% improvement in edit correctness on our test set.
copyright: none
@@ 1. Introduction
While the first major win for the use of Large Language Models (LLMs) in coding came from autocomplete (cop, 2025; Ziegler et al., 2022), where the cursor position and the surrounding code were the main inputs to an LLM, the emphasis has shifted to situations in which some amount of developer input – or prompt — is expected to convey the intent to the LLM. This holds, whether it is for code generation, or editing/transformation of existing code, or indeed as a part of a multi-step interaction between the programmer and the LLM.
A notable difference between the autocomplete and prompt-based features is the need for developers to write the prompts. To receive the desired output from such tools, developers need to clearly elaborate their needs, which can require both time and cognitive effort. As much previous research in understanding professional developers’ perceptions and productivity gains was conducted with autocomplete as the main focus (Peng et al., 2023; Paradis et al., 2024; Ziegler et al., 2022), we still have unanswered questions on developers’ prompting strategies for code editing tools.
Transform Code. We specifically focus on the case of explicitly prompted code editing or transformation, and conduct a case study of Transform Code (Figure 1). Transform Code is a feature in internal IDE at Google, powered by an LLM trained on many popular programming languages, wherein developers can modify and extend code with custom natural language prompts in context of a selection of code within the editor. When a developer selects a certain range of code and activates the feature using a shortcut or by clicking a button, a small Transform Code window pops up where the developer can enter a short natural language prompt indicating the edit or transformation that they would like to see applied to that code. Once the LLM provides a response, the suggested code edit is displayed as a diff in the original code, and the developer can either accept or reject it.
For instance, Figure 1 shows a request by a developer to “simplify” a piece of code (a), who and receives a suggested code edit that simplifies the highlighted code (b). However, often times, the suggested code edits developers receive can be different from what they expected. In Figure 1 for example, the suggested code edit changes the functionality to simplify it, which is likely not the developer’s intention.
The above use of prompts, however, is different from that in prompt-driven code generation that is now popular with modern LLMs (Li et al., 2022; Roziere et al., 2023; Li et al., 2023); the specific use of prompts here is for in-IDE local code transformations. The prompts are typically short, a single sentence of around 15 words, but typically even shorter.
@@
(b) Once the LLM provides a response, the code edit is displayed as diff in the developer’s original code. The developer can then either Accept or Discard the suggested code edits (using the first two buttons from the left), or retry the prompt (third button from left).
@@
We aim to understand the prompting behavior of developers when using Transform Code through a multi-phased study of Transform Code logs. Do prompts generally work as expected? If not, how do people react to unsuccessful prompts. For instance, do people try the same prompt multiple times? Do they alter their prompts and retry? Finally, how might we assist developers in providing prompts that will produce suggested code edits that align better with their expectations?

### v2

# Understanding and supporting how developers prompt for LLM-powered code editing in practice

@@ Abstract.
Large Language Models (LLMs) are rapidly transforming software engineering, with coding assistants embedded in an IDE becoming increasingly prevalent. While research has focused on improving the tools and understanding developer perceptions, a critical gap exists in understanding how developers actually use these tools in their daily workflows, and, crucially, where they struggle. This paper addresses part of this gap through a multi-phased investigation of developer interactions with an LLM-powered code editing feature, Transform Code, in an IDE widely used at Google. First, we analyze telemetry logs of the feature usage, revealing that frequent re-prompting can be an indicator of developer struggles with using Transform Code. Second, we conduct a qualitative analysis of unsatisfactory requests, identifying five key categories of information often missing from developer prompts. Finally, based on these findings, we propose and evaluate a tool, AutoPrompter, for automatically improving prompts by inferring missing information from the surrounding code context, leading to a 27% improvement in edit correctness on our test set.
conference: 48th International Conference on Software Engineering; April 2026; Rio de Janeiro, Brazilcopyright: none
@@ 1. Introduction
(b) Once the LLM provides a response, the code edit is displayed as diff in the developer’s original code. The developer can then either Accept or Discard the suggested code edits (using the first two buttons from the left), or retry the prompt (third button from the left).
@@
While the first major win for the use of Large Language Models (LLMs) in coding came from autocomplete (copilot; ziegler2022productivity), where the cursor position and the surrounding code were the main inputs to an LLM, the emphasis has shifted to situations in which some amount of developer input – or prompt — is expected to convey the intent to the LLM. This holds whether it is for code generation, editing existing code, or as part of a multi-step interaction between the programmer and the LLM.
A notable difference between the autocomplete and prompt-based features is the need for developers to write the prompts. To receive the desired output from such tools, developers need to clearly elaborate their needs, which can require both time and cognitive effort. As much previous research on understanding professional developers’ perceptions and productivity gains has been conducted with autocomplete as the main focus (peng2023; paradis2024how; ziegler2022productivity), we still have unanswered questions on how developers prompt in practice.
Transform Code. We specifically focus on the case of explicitly prompted code editing or transformation, and conduct a case study of Transform Code (Figure 1). Transform Code is a feature in the internal IDE at Google. Developers can modify and extend code with custom natural language prompts in context of a selection of code within the editor, similar to features like GitHub Copilot’s Edit Mode (copilot) or Cursor’s Inline Edit (cursor). It is powered by an LLM trained on many popular programming languages, which is comparable to models like Codex (chen2021evaluating).
When a developer selects a certain range of code and activates the feature using a shortcut or by clicking a button, a small Transform Code window pops up where the developer can enter a short natural language prompt indicating the edit that they would like to see applied to that code. Once the LLM provides a response, the suggested code edit is displayed as a diff in the original code, and the developer can either accept or reject it. For instance, Figure 1 shows a request by a developer to “simplify” a piece of code (a), and receives a suggested code edit that simplifies the highlighted code (b). By providing in-situ support, Transform Code provides easy contextualization and minimizes context switching.
However, while considerable research has been conducted on prompt engineering for code generation (chen2025empirical; errica2024did), mainly for benchmarks such as HumanEval (chen2021evaluating), research on the prompting behavior of developers in the enterprise setting remains sparse. For example, we observed that developers typically use short, simple prompts mostly with a single sentence of around 15 words. This shows the actual developer prompting patterns are distinct from the code generation benchmark settings, where many prompting engineering approaches utilize a full description of functions with additional information (e.g., test cases (mathews2024test)),
In this paper, we address this gap by investigating how developers interact with LLM-powered code editing tools in practice, how they struggle, and a method for mitigating their struggles through a multi-method, multi-phase investigation. In Phase I (Section 3), we carry out, to our knowledge, the first log-based study of developers’ behavior when it comes to prompting for code editing. With extensive fine-grained logs, we analyze how developers interact with Transform Code in their daily workflow, with enterprise-level tasks and complexity. This log-based study minimizes many biases (dell2012yours) that typically exist with studies asking developers how they feel. We find that around 7% of the Transform Code requests are identical re-prompting requests or only minor editing, potentially signaling their struggle in getting a satisfactory response from Transform Code.
To understand the reasons behind these developers’ struggles and propose a mitigation, we build a more focused dataset of unsatisfactory Transform Code requests (Section 4). In addition to the original user prompt and the Transform Code response, we approximate the “ground truth” answer expected by the developer by looking at a future code snapshot or developer-submitted feedback. This allows us to ask, counterfactually, what would have been a prompt that produced that ground truth answer, or something very close to it.
Using this data, in Phase II (Section 5), we conduct a qualitative study to understand the common issues with the original developer prompts that prevented Transform Code to generate the ground truth code edit. We find five common types of information missing from prompts: specifics, operationalization plan, localization/scope, codebase context, and developer intent.
Finally, in Phase III (Section 6), we test the feasibility of automatically enhancing the user prompts to mitigate the user struggles. We present a prototype tool, AutoPrompter, that considers the common issues of user prompts identified, and automatically augments the user prompts with relevant contextual information. We find that this tool can improve the effectiveness of developer prompts by around 27%, automatically and without further developer input.
2. Related work
2.1. AI Developer Tools for Professional Developers
In recent years, driven by significant improvements in large language models (LLMs), many developer tools have been built upon or have incorporated LLMs. GitHub Copilot (copilot) is one of the earliest LLM-powered developer tools, suggesting code in real time, based on context. More similar tools have been built, such as Cursor (cursor), Windsurf (windsurf), and Bolt (bolt), supporting code completion, editing, and generation, or even agent-based support. Significant effort has also been invested in AI-enhanced software development tools for internal use, such as at Meta (bader2021) and Google (AIsoftEng; MLcodecomplete; froemmgen2024), to support their developers’ workflows and proprietary codebases. Many have conducted studies to understand their potential to enhance developer productivity, investigating both perceived (bird2023; murillo2023) and actual (peng2023; paradis2024how) productivity gains, through various methodologies. Research has also studied challenges developers face when using AI tools. A notable struggle reported multiple times is in effectively communicating requirements to AI tools, which often requires developer cognitive effort (parameswaran2024revisiting; nam2024understanding; vaithilingam2022; liu2023what; tafreshipour2024prompting).
Analyzing telemetry logs to understand usage patterns (murphyhill2018discovering; nam2024understanding) and improve features (bibaev2022all; froemmgen2024) is a common practice in software engineering. However, AI tools have only recently been broadly adopted by developers for real-world usage and are rapidly evolving. Consequently, research on how developers incorporate these tools into their actual workflows has been limited. More recently, some efforts have been made to measure the overall impact of AI tools at scale, such as by measuring acceptance rates (ziegler2022productivity) or the number of pull requests (peng2023). However, these studies primarily focused on quantifying productivity gains rather than understanding how developers use the tools. One notable exception is mozannar2024, which collected and analyzed telemetry data from programmers using Copilot. While that paper provides valuable insights into programmer activities using telemetry segments, the data was collected from only 21 participants and the analysis was done after manual labeling. In contrast, our analysis examines logs from developers at Google, reporting insights at a large scale.
2.2. Approaches for Prompt Engineering and Enhancement
As early research has identified challenges in writing effective prompts (parameswaran2024revisiting; nam2024understanding; vaithilingam2022; liu2023what), researchers and practitioners have explored ways to assist developers in prompt writing. Several (semi-) automatic approaches have been proposed, e.g., meta-prompting (yang2024large; fernando2024promptbreeder; zhou2023large) for automatically optimizing prompts in a systematic fashion. As another example, it was demonstrated in Automatic Prompt Engineer (APE) (zhou2023large) that LLMs can discover human-level prompts through a search process, outperforming manually written prompts on multiple benchmarks. However, the main research efforts have focused on optimizing the system instruction prompt that can be used for several distinct tasks within the same category, often requiring many examples to automatically quantify the quality gains from prompt improvement. Another notable example is prompt augmentation tools for image generation, like LMOps (hao2023optimizing), which aim at generating diverse images through augmented prompts. However, their main goal is to generate various outputs, rather than satisfying more constraints by inferring missing specifications from the surrounding context, which is the goal of prompt enhancement for developer tools.
In software engineering domains, researchers have proposed various strategies to mitigate challenges in prompting. Several attempts have been made to enhance code generation quality through techniques such as adding test cases (mathews2024test), planning mechanisms (jiang2024self), or leveraging semantic entropy distributions of generated programs (jia2025automated). However, these approaches primarily focus on optimizing code generation performance by encouraging LLMs to engage in more sophisticated reasoning, rather than supporting the diverse code editing tasks that developers encounter in practice. Other researchers have proposed proactively providing assistance to reduce the cognitive burden on developers during prompt engineering (nam2024understanding; need2025chen). These approaches are useful when users are unfamiliar with a domain or when their intent is easily inferred. However, it is unclear whether those benefits can be generalized in enterprise environments, where developers work on complex, specialized tasks that make intent inference substantially more difficult. Some studies have explored enhancing initial prompts by requesting clarification questions (li2022python), which can be highly effective for correctly inferring user intent. However, this approach imposes cognitive overhead on users, who must read follow-up questions and provide appropriate responses before receiving the desired output. To the best of our knowledge, ours represents the first study to analyze the prompt strategies employed by professional developers in their daily workflows, while investigating the feasibility of directly enhancing user prompts to minimize cognitive load from multi-turn clarification or debugging processes.
3. Phase 1: Transform Code Log Analysis
Transform Code has been adopted by a significant fraction of developers since its release. Through iterative improvements to both the underlying LLM and the user interface, Transform Code has become a core part of developers’ workflows. Despite its positive reception and appreciation, generating code edits from natural language prompts remains a relatively new concept for developers. Thus, some developers have expressed challenges, mirroring the experiences of users of similar tools (kruse2024can; nam2024understanding).
In the ideal case, a developer using Transform Code would send a request with a prompt for a given task, Transform Code would provide an appropriate code edit, the developer would review the edit, and proceed to the next task. However, in practice, it is unrealistic to expect an LLM-powered tool to provide the perfect solution on the first attempt. In this section, we analyze telemetry logs from Transform Code and the IDE, focusing on how developers use Transform Code, and how they react when Transform Code does not provide the ideal code edit they anticipated.
3.1. Data
3.1.1. In-IDE Event Logs
3.1.2. Transform Code Logs
3.1.3. Dataset used for analysis
From this combined dataset, we sampled data of over 10,000 individual developers who used Transform Code at least 24 times during the month, which is the median usage of the entire population. This threshold was set to minimize noise arising from non-regular usage of Transform Code, e.g., where new users might test Transform Code with simple prompts like ”hi.”
3.1.4. Limitations
(a) After rejection.
(b) After acceptance.
3.2. Post-Transform Code Interactions
Figure 2 visualizes the two consecutive in-IDE developer actions after requesting Transform Code. We first analyzed the actions made after rejection, Figure 2-(a), to understand how developers cope with unsatisfactory Transform Code edits. Then, we also analyzed the actions made after acceptance, Figure 2-(b), to check for any differences in developers’ reactions. This distinction is also important because the Transform Code edit is applied to the user’s code upon acceptance, but discarded upon rejection. Thus, subsequent edits made by the user in the same parts of the code after acceptance likely indicate refinement or adjustment of the applied Transform Code edit, whereas actions taken after rejection indicate the user is continuing to work on the original code.
3.2.1. Post-rejection
The most common subsequent action after rejecting a Transform Code edit was Type (33.5%), including typing a character or inserting a line. This indicates that developers often start working on their tasks by themselves after checking an unsatisfactory Transform Code edit. A particularly interesting next action was a repeated request to Transform Code (24.3%), which is the second most common action after rejecting a Transform Code edit. This likely indicates that developers re-prompted Transform Code for the same task. We found that not an insignificant proportion of developers (22.2%) made even another Transform Code request, possibly indicating developer struggle in using Transform Code, and/or their desire to delegate the task to Transform Code instead of picking up on the task themselves, like those who started typing. The following frequent subsequent actions were Delete (20.1%), Paste (9.2%), and Cut (3.1%), showing other mechanisms of editing the original code, possibly adopting code from code or documentation search.
3.2.2. Post-acceptance
The most common subsequent action was Delete (32.9%), and the second most frequent action was Type (28.7%). This indicates that the developers frequently make edits to the Transform Code-edited code, after accepting a response, to further refine it. The next most frequent action was a repeated request to Transform Code (11.9%). Even after accepting the Transform Code edit, users requested Transform Code from the same code region more than 10% of the time. This could indicate that developers found unsatisfactory edits after accepting the edits, and re-prompted for the same task, or it is still possible that developers moved on to the next task and requested another code edit within the same part (e.g., adding documentation after implementing a function). Developers also made edits via other mechanisms, possibly to the Transform Code-edited code, by pasting code (5.8%), or cutting the code (5.1%).
3.3. Characteristics of Re-prompting
3.3.1. Frequency of Identical Re-prompting
We first investigated how often developers re-prompt with an identical prompt. We started with identical re-prompting, as it is very likely that the user is making a follow-up request, instead of moving on to the next task after getting a satisfactory response.
3.3.2. Re-prompting with tweaks
3.4. Discussion
Overall, we observed that developers often do not invest sufficient time in crafting effective prompts when using Transform Code. Although documentation and the Transform Code user interface instructed users to provide specific details and necessary context, we found that few users did so when tweaking the prompt. Developers typically resorted to manual editing or invested minimal effort in refining the prompt for a subsequent attempt when an initial, low-effort attempt with Transform Code yielded unsatisfactory code edits. This behavior suggests that there may be an ‘economics of prompting,’ where developers weigh the costs of prompting against the expected benefits. Thus, to improve the user experience and maximize the benefits of Transform Code, enhancing the underlying model performance alone is not sufficient; efforts must also focus on reducing users’ cognitive load during prompting and on helping them write prompts more effectively (errica2024did; nam2024understanding; liu2023what).
4. Unsatisfactory Code Editing Dataset
To inform the mitigation strategies that help developers write more effective prompts for Transform Code, we aimed to understand current prompting practices and identify areas for improvement. We collected a dataset of real-world Transform Code usage, specifically focusing on developer requests that yielded not satisfactory outcomes.
4.1. User Submitted Unsatisfactory Usage
Transform Code has a feedback mechanism, where developers unhappy with the code edit suggestion from Transform Code can submit their desired code edit instead. This improved suggestion is bundled with the original prompt, code context, and original Transform Code suggested code edit upon submission. For these examples, the desired code edits most probably satisfy the user intent and expectations, as developers themselves need to deliberately create and submit the improved rewrite. However, not all developers are motivated to invest time in submitting such rewrites; therefore, the number of these examples is quite limited. After excluding the examples used for the fine-tuning of the Transform Code model, we had 52 examples left for this analysis.
4.2. Automatically Extracted Unsatisfactory usage
To obtain the desired code edit, we extracted a future snapshot of the code, assuming that regardless of how (un)satisfied the developers were, the task would have been resolved at some later point in time. We consider the code state after 30 minutes, to leave enough time for developers to make their desired edits, especially for cases where the task requires finding additional information. However, given that the edit time can vary significantly, depending on the task complexity and the user expertise, it was possible that code edits not relevant to the original Transform Code request could have been included. Developers could also potentially give up on finishing the task, or work on different tasks in the meantime. This made manual verification necessary.
  • the desired code edit only removes code from the original code.
4.3. Full Dataset for Analysis and Solution Building
Combining the data collected from the two sources resulted in 159 examples. Among these, we used 111 examples (70%) of the dataset for understanding developer prompting practices and issues, and we set aside 48 (30%) to be used for testing the solution we would build based on our gained understanding. Among these, we excluded examples that request code documentation only or simple renaming, where it is challenging to understand why the original Transform Code code edits were not satisfactory, and there is a subjective component involved (see Table E-Example 3 in (artifact)). This resulted in a training dataset of 72 examples, and a test dataset of 33 examples. As the data were collected from real-world logs containing proprietary code, we cannot share the entire dataset, but we can only report the results and share examples.
5. Phase 2: Understanding Common Issues of Developer Prompts
With the collected dataset, we conducted an error analysis of the developer requests that resulted in unsatisfactory code edits from Transform Code, to understand what would be required in a better prompt. We leveraged an LLM111in our case, Gemini, but we could have used any of the popular LLMs. to help us identify for each example the missing ingredients in a prompt, given prior knowledge of the complete code context, the prompt, and the desired edit. This is based on the premise that with such a better prompt, Transform Code would likely have generated the desired solution.
5.1. Error Analysis using an LLM
Code # Description
Specifics 28 Missing or under-specified specifications, for example, the prompt lacks value information
Operationalization plan 28 Missing or under-specified information on what operations need to be performed or what functions need to be used, or what edge case handling is needed to get to the desired code edit
Codebase context 11 Prompt itself did not provide the necessary background information for the model to be able to generate the desired code edits, including missing information about the broader system, existing codebase. Developers may have assumed that the model had all the necessary context.
User intent 12 Missing or under-specified user intent in the prompt; it is not clear what the developer’s goal is or what type of code edit they want to achieve
To collect the common issues with user prompts (i.e., what is missing from the prompt so that Transform Code gives the desired result), we presented the LLM with: (1) the original user prompt, (2) the code context, (3) the original Transform Code edit, and (4) the desired code edit. Then we asked the LLM to produce (a) an explanation of what was missing from the original user prompt and (b) an improved prompt. See Appendix A (artifact) for the full system prompt used.
5.1.1. Feasibility Testing
As a sanity check to test whether an LLM can reason about the issues in the user prompts, we tested whether Transform Code can replicate the desired edits from the LLM-enhanced prompts, when provided with the desired edits to enhance the prompt. Out of 72 examples, we found that 43 enhanced prompts (59.72%) enabled Transform Code to replicate the desired edits. These numbers indicate that the LLM is able to pinpoint the issues with the user prompts well, and with all the provided information in the prompt, Transform Code is able to generate satisfactory outcomes for a fairly large number of cases. At the same time, we found that not every unsatisfactory Transform Code edit is due to an incomplete user prompt. For example, cases that cannot be easily abstracted into a concise natural language prompt (e.g., adding tens of import files without clear patterns across them) could not be improved by rewriting prompts even after seeing the desired edit. We also discovered that in some examples, Transform Code provided better code edits than the desired code edits. Here the developer may have accepted the Transform Code suggestion, had they seen it, but as we do not know the developer’s original intent, they were not considered semantically identical. Finally, the underlying model has limitations; it may not be able to generate the desired code edit even when presented with the perfect prompt.
5.2. Qualitative Analysis
We qualitatively analyzed the issues of user prompts for Transform Code extracted from the 72 examples in our dataset, using thematic analysis (braun2006using). Our qualitative analysis consisted of several phases. Three authors analyzed the original user prompts and the feedback provided by the LLM, which indicated what information was missing from the original prompt, using inductive coding (fereday2006demonstrating). As we reviewed the data, we generated an initial set of categories to describe the different types of missing context and specifications. If we could not clearly identify which information was missing from the original prompt, we referred to the code context, the original Transform Code edit, and the desired code edit, which also helped mitigate any errors coming from potential hallucinations. We also wrote analytic memos during the discussion, including observations of patterns and potential connections among the categories.
5.3. Results
The analysis resulted in 5 codes of missing or under-specified information in user prompts for Transform Code. Table 1 shows the codebook (ando2014achieving), including the code, the frequency of the codes, and the description. Each example could be labeled with up to 2 codes.
Codebase Context. Codebase context refers to cases where the user assumes that Transform Code would know all the background knowledge and details about the codebase they are working on, in cases where Transform Code actually had no access to such information. This could happen because end users might not have a good understanding of how the underlying model is trained, which data are used, and what additional context is included as part of the request, which may have led the developers to assume a level of contextual awareness that the model does not possess. For example, the desired code edit that Transform Code was not able to produce given “add [library]” was successfully replicated with an improved prompt, “Add “[path to the library]” as a dependency to the [build] rule“.
User Intent. A frequent example of this category is a user asking Transform Code to perform a rather vague and open-ended task. For example, for a request with the original user prompt “make it simpler”, the LLM pointed out that “The prompt did not specify the intended behavior when [variable1] is present in [variable2]. The model incorrectly interpreted “make it simpler” as removing the redundant assignment instead of preserving the original conditional logic with the necessary ‘continue’ statement.” The improved prompt generated by the LLM, “Refactor the conditional logic for assigning [variable3] to maintain the original behavior, adding a ‘continue’ statement when [variable1] is in [variable2].”, successfully replicated the desired code edit. When prompts lack user intent, compared to other issues, it becomes extremely difficult for Transform Code to produce the desired code edit, as it almost requires reading the minds of developers. Thus, if the desired code edit matches the most likely code edit to be made in general, given the code context, Transform Code can produce a good code edit, but if not, it is very challenging to achieve the desired code edit.
Satisfactory Prompts. We identified five examples from the dataset where the original user prompts appeared satisfactory without requiring further improvement, at least from the researchers’ perspective. Because the dataset is a collection of user requests that resulted in unsatisfactory code edits, rather than a collection of inadequate prompts, it is reasonable that some of these unsatisfactory code edits are not attributable to the user prompts.
5.4. Feasibility of Automatic Analysis
In this study, we conducted a qualitative analysis following the traditional thematic analysis process, as this was the first attempt to understand user prompt issues for LLMs in the context of code editing. However, this process of extracting insights from usage data could potentially be automated with LLMs, similar to other attempts in conducting qualitative research with LLMs (bano2023exploring; meng2024exploring), particularly to reduce researcher effort and cost. To assess the feasibility of such automation, we instructed an LLM to carry out the open coding and codebook construction. We found the LLM-generated codebook to be largely similar to the one created by researchers (see Appendix C (artifact)). This result suggests the potential for applying automated analysis with low cost to understand user challenges in other tool contexts (e.g., automatic program repair (rondon2025evaluating)) or domains (e.g., code migration (nikolov202how)).
6. Phase 3: Automatic User Prompt Enhancement
In the previous section, we identified categories of common issues with user prompts. In this section, we tested the feasibility of automatically improving user prompts by exploiting this finding, and tested whether some of the missing or underspecified user requirements could be inferred from the code context.